OpenStudy (kaptain_mittens):

MEDAL AND FAN PLEASE HELP ME!! Evaluate. 5^square root of 8 times square root of 8 divided by 3^square root of 8 to the power of 5

OpenStudy (p0sitr0n):

something like 5*root8 * root8 / 3*root8^5 ?

OpenStudy (kaptain_mittens):

no the 5 is like the cube root of square root of 8 and the 3 is the cube root of the square root of to the power of 5, i think its called cube root. Its like an exponent on the opposite side of the number

OpenStudy (kaptain_mittens):

Here

OpenStudy (kaptain_mittens):

It's late so I need to finish asap

OpenStudy (kaptain_mittens):

P0sitr0n?

OpenStudy (anonymous):

still need help

OpenStudy (kaptain_mittens):

yes!

jimthompson5910 (jim_thompson5910):

are you able to convert each radical to exponent form?

OpenStudy (kaptain_mittens):

Well I know that the 8 on top becomes 8 to the power of 1/5 but after that I dont know what to do

jimthompson5910 (jim_thompson5910):

I think you mean \[\Large \sqrt[5]{8} = 8^{1/5}\] right?

OpenStudy (kaptain_mittens):

yes

jimthompson5910 (jim_thompson5910):

how about the square root of 8?

OpenStudy (kaptain_mittens):

I don't how to simplify it from that point

OpenStudy (kaptain_mittens):

which is pretty sad i know

jimthompson5910 (jim_thompson5910):

no, it's not

jimthompson5910 (jim_thompson5910):

\[\Large \sqrt{8} = 8^{1/2}\]

jimthompson5910 (jim_thompson5910):

then how about the cube root?

OpenStudy (kaptain_mittens):

where did the 2 come from

jimthompson5910 (jim_thompson5910):

it turns out that square roots are simply radicals with index of 2 \[\Large \sqrt{8} = \sqrt[2]{8}\]

OpenStudy (kaptain_mittens):

oh ok

jimthompson5910 (jim_thompson5910):

but you don't need to write that little "2" because it's understood that square roots are radicals with index 2 it's like x/1 = x in a way

jimthompson5910 (jim_thompson5910):

how about the cube root?

OpenStudy (kaptain_mittens):

it becomes 8^3/5?

jimthompson5910 (jim_thompson5910):

oh wait, no, I mixed up the two

OpenStudy (kaptain_mittens):

so right now we have 8^1/5 and 8^1/2 over 8^3/5

OpenStudy (kaptain_mittens):

what?

jimthompson5910 (jim_thompson5910):

it should be \[\Large \sqrt[3]{8^5} = 8^{5/3}\]

jimthompson5910 (jim_thompson5910):

so we go from this \[\Large \frac{\sqrt[5]{8}\sqrt{8}}{\sqrt[3]{8^5}}\] to this \[\Large \frac{8^{1/5}*8^{1/2}}{8^{5/3}}\]

OpenStudy (kaptain_mittens):

alright so the top simplifies to 64^7/10

OpenStudy (kaptain_mittens):

or is it 64^1/10?

jimthompson5910 (jim_thompson5910):

no you don't say 8*8 = 64 in the numerator the base stays 8 the whole time

OpenStudy (kaptain_mittens):

oh

OpenStudy (kaptain_mittens):

so its 8^7/10

jimthompson5910 (jim_thompson5910):

yeah up top

OpenStudy (anonymous):

bye

OpenStudy (kaptain_mittens):

ok so what do we do from here

jimthompson5910 (jim_thompson5910):

we have this now